OpenStudy (anonymous):

would 4 radical 2 + 9 radical 3 be the answer to: radical 2 + 3 radical 3 + 3 radical 2 + radical 27 if not what is and how did you find it

OpenStudy (amistre64):

wolframalpha.com is pretty fast

OpenStudy (amistre64):

and it would be helpful if i clicked on the right posts that im responding to ...

OpenStudy (amistre64):

consider the radicals of like numbers to be like terms

OpenStudy (amistre64):

3+3 = 6 not 9

OpenStudy (anonymous):

o so it would be 4 radical 2 + 6 radical 3 not 9 radical 3 correct??

OpenStudy (amistre64):

\[\sqrt{2}+3\sqrt{2}+3\sqrt{3}+\sqrt{27}\] \[\sqrt{2}+3\sqrt{2}+3\sqrt{3}+\sqrt{9*3}\] \[\sqrt{2}+3\sqrt{2}+3\sqrt{3}+3\sqrt{3}\] \[(1+3)\sqrt{2}+(3+3)\sqrt{3}\]

OpenStudy (amistre64):

4 and 6 yes

OpenStudy (anonymous):

thank you